Baldwin Company's product Best has Fixed Costs of $12,000, sales of 10,000 units, selling price per unit of $30, and a contribution margin ratio of 40%. Baldwin is considering investing in a Marketing program that costs $20,000 and is projected to increase sales volume by 10%. All else constant, if Baldwin makes this decision based on a minimum acceptable ROI in the first year of 50% for the project, should they invest in this marketing program?  

The answer is "Yes, because the investment will increase Baldwin's Contribution Margin by $12,000 and their Fixed Costs by $20,000 resulting in a 60% ROI for this investment in the first year "

But I don't understand how the contribution margin increase by $12,000.
Answer #1

Earlier the sales was 10,000 units @ $30.

So, the sales were 10000×30 = $3,00,000.

And the contribution margin is 40%.

So, contribution = 3,00,000×40% = $1,20,000.

Now, the sales is increasing by 10%.

So, the sales is 3,00,000×110% = $3,30,000.

Only fixed cost increases by 20,000. Else every thing remains constant.

So, contribution margin is 3,30,000×40% = $1,32,000.

In this way, contribution margin increase by $12,000.